OpenStudy (anonymous):

Suppose that the number of bacteria in a certain population increases according to an exponential growth model. A sample of bacteria selected from this population reached the size of bacteria in three hours. Find the continuous growth rate per hour. Write your answer as a percentage. Do not round any intermediate computations, and round your percentage to the nearest hundredth.

OpenStudy (amistre64):

P = P(%) e^(3r) maybe

OpenStudy (amistre64):

whole = sample e^(3r) whole ----- = e^(3r) sample ln(w/s) ------ = r ; might want to throw on a constant of variation ??? 3

OpenStudy (amistre64):

this is hard to parse: "A sample of bacteria selected from this population reached the size of bacteria in.."